OpenStudy (anonymous):

Prove \[\forall a \in \mathbb{R}, a * 0 = 0\]

OpenStudy (anonymous):

use the fact that 0 + 0 = 0

OpenStudy (anonymous):

you are trying to prove that for all real constants a, a*0 = 0 cant you just say that you multiply a*b you add b a tmes. thus a*0 is 0+0+0+0 a times and as 0+0=0 a*0 is also 0?

OpenStudy (anonymous):

or is not mathematical enough

OpenStudy (anonymous):

that's what I was trying to prove: a*0 = 0. your proof assumes that it is already true, when I only know for a fact that 0 + 0 = 0, but I don't know yet that a*0 = 0

OpenStudy (anonymous):

sumbul to the rescue :-D

OpenStudy (anonymous):

consider ax0=ax(1+(-1)) (by existence of additve inverse) =aX(1)+aX(-1) by distributive law =a+a(-1) by existence of multiplicative identity =a+(-a) (because a(-1)=-a) =0 by existence of additve identity

OpenStudy (anonymous):

that works :-D

OpenStudy (anonymous):

yes and it is a valid prove we did it in real analysis